OpenStudy (anonymous):

3.7(-83x-3)=3.7(7x-30) Help Guys!

OpenStudy (anonymous):

Cancel 3.7 on both sides Then bring all constants to one side and all variables to another side. You would get 3/10 as the answer

OpenStudy (yanasidlinskiy):

Do you know the steps on solving this situation??

OpenStudy (anonymous):

Simplifying 3.7(-83x + -3) = 3.7(7x + -30) Reorder the terms: 3.7(-3 + -83x) = 3.7(7x + -30) (-3 * 3.7 + -83x * 3.7) = 3.7(7x + -30) (-11.1 + -307.1x) = 3.7(7x + -30) Reorder the terms: -11.1 + -307.1x = 3.7(-30 + 7x) -11.1 + -307.1x = (-30 * 3.7 + 7x * 3.7) -11.1 + -307.1x = (-111 + 25.9x)

OpenStudy (anonymous):

-11.1 + -307.1x = -111 + 25.9x

OpenStudy (anonymous):

Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right.

OpenStudy (anonymous):

Add '-25.9x' to each side of the equation. -11.1 + -307.1x + -25.9x = -111 + 25.9x + -25.9x

OpenStudy (anonymous):

-307.1x + -25.9x = -333x -11.1 + -333x = -111 + 25.9x + -25.9x

OpenStudy (anonymous):

25.9x + -25.9x = 0.0 -11.1 + -333x = -111 + 0.0 -11.1 + -333x = -111

OpenStudy (anonymous):

Add '11.1' to each side of the equation. -11.1 + 11.1 + -333x = -111 + 11.1

OpenStudy (anonymous):

-11.1 + 11.1 = 0.0 0.0 + -333x = -111 + 11.1 -333x = -111 + 11.1

OpenStudy (anonymous):

-111 + 11.1 = -99.9 -333x = -99.9

OpenStudy (anonymous):

Divide each side by '-333'.

OpenStudy (anonymous):

x = 0.3

OpenStudy (anonymous):

THANK YOUUUU GUYS <3

OpenStudy (anonymous):

welcome @takisloverr